# Towards Realistic Data Generation for Real-World Super-Resolution

### v3

@@ 3.1 Experiments Settings
Real-ESRGAN [33] 24.3974 0.6798 0.3881 157.5237 0.2618 0.7552
BSRGAN [32] 23.2665 0.6043 0.5093 177.9996 0.3048 0.6783
SynDiff [46] 25.2461 0.7588 0.2371 119.9950 0.1944 0.7984
RealSR [28] Ours 26.1615 0.7940 0.2226 107.2375 0.1865 0.8238
RealESRGAN [33] 26.3510 0.6935 0.3842 60.2665 0.2437 0.7710
BSRGAN [32] 26.1970 0.6869 0.4319 62.9189 0.2646 0.7137
SynDiff [46] 28.0125 0.8136 0.1739 31.0855 0.1673 0.8462
DRealSR [31] Ours 28.6629 0.8360 0.1497 30.0078 0.1567 0.8577
RealESRGAN [33] 26.0680 0.5981 0.5202 106.3958 0.2855 0.7231
BSRGAN [32] 27.6215 0.7179 0.4399 105.3289 0.2760 0.6771
SynDiff [46] 28.5020 0.8075 0.2423 70.6379 0.2044 0.8213
@@
3.2 Quantitative Results
Generalization Ability of Real SR Models. We compare our method with the existing data generation methods on both PSNR-oriented and Perceptual-oriented Real SR models to validate the superiority of our method in boosting the generalization capabilities for real-world image super-resolution, and the results are shown in Tab. 1. We can observe that our method comprehensively improves the performance of PSNR-oriented and Perceptual-oriented SR models across three benchmarks. It’s worth noting that our approach achieves a significant performance improvement, including 0.75 dB in RRDB [54] on the PSNR of the RealSR benchmark; 0.296dB and 0.699 dB in SwinIR [55] and light-weight SwinIR-L on the PSNR of the RealSR benchmark. Furthermore, our methods also significantly improve LPIPS and FID in four SR models, including 0.0395 and 12.619 in HAT [56] on the LPIPS and FID of RealSR benchmark. This demonstrates that our method effectively improves the generalization capacity of popular SR models.
Accuracy of Generated Real LR. To validate our method’s superiority in generating accurately realistic and matched real LR, we conduct comparisons with existing methods in terms of the accuracy of the generated data on six metrics for evaluation. Specifically, we utilize HR images in the three real-world datasets as the content reference and employ the existing methods to synthesize generated LR images, while our method is further able to utilize the degradation reference from real LR. As shown in Tab. 2, we can observe that our method achieves the best performance on three datasets under six metrics, including distribution distance metric: FID, image structure and texture similarity metric: DISTS, and perceptual metric: LPIPS and CLIP-score, etc. It demonstrates the superiority of our Decoupled DDPM and verifies that our generation method is closer to target real LR images.
Generalization Ability on Out-of-distribution Data. Although we have collected about 152,000 real LR images for training, unseen scenarios often arise in real-world applications. To further explore our generalization capabilities on out-of-distribution data, we randomly extract 200 images from the unseen real-world video SR data [64] for evaluation, denoted as RealVSR. Then, we capture the degradation representing those images to generate the training data and compare it with existing methods on the SwinIR-L model, and the results are shown in Tab. 5. We can observe that our method still achieves the best performance on the out-of-distribution RealVSR benchmark, surpassing the existing state-of-the-art methods by 0.2214 and 0.0055 in PSNR and SSIM, respectively. This demonstrates the ability of our method to generalize for unseen scenarios in the real world.
@@ 3.3 Qualitative Results
We visualize the results of the generated LR images, as shown in Figure 3. It is clear that the Real-ESRGAN and BSRGAN, which are unable to perceive real degradation, exhibit a significant gap compared to real LR images. The SynDiff also struggles to adaptively capture the real degradation and content representation, resulting in unreal degradation distribution and low fidelity, as shown in the third row in Figure 3. However, benefiting from our adaptive perception ability of the degradation of real LR image and content of HR image, our RealDGen achieves the best visual effect in terms of both the realism of degradation and the fidelity of content. Furthermore, to demonstrate the visual effects of Real SR model, we conduct visualizations of the SwinIR model in the RealSR and DRealSR benchmarks, as shown in Figure 4. It is clear that models based on SynDiff, BSRGAN, and Real-ESRGAN exhibit artifacts and blur. However, our method achieves the best visual quality without artifacts. The textual scene in the second row of Figure 4 also demonstrates the superiority of our method for text texture. This showcases the superior visual effects of our method in real-world scenarios. More visual comparison results are presented in Section A.7 of the appendix.
@@ 3.4 Ablation Study
Degradation and Content Extractor. Our core idea is utilizing the degradation extractor $E_{deg}$ and content extractor $E_{cont}$ to extract degradation and content representation to control Decoupled DDPM in generating realistic LR. To verify this, we eliminate components $E_{deg}$ and $E_{cont}$, subsequently training the decoupled DDPM and fine-tuning the remaining extractor. For quick evaluation, we employ SwinIR-L as the baseline and evaluate on the Smartphone benchmark. The results in Tab. 4 demonstrate that the absence of $E_{deg}$ and $E_{cont}$ will cause mismatch and unreal degradation and content distortion problems and result in performance degradation, thereby affirming the indispensability of $E_{deg}$ and $E_{cont}$ in our method.
$T$ in Data Generation. As illustrated in Sec. 2.3, in inferencing, we denoise from an initial LR image $\mathbf{x}_{t}$ with $t$ step noise added and perform t steps of denoising to improve the fidelity of Decouple DDPM, where $t$ is selected from the range of 0 to $T$. To explore it, we set $T$ to 200, 300, 400, and 500 and conduct ablation experiments, as presented in Tab. 4. It is observed that the optimal $T$ is 300. A lower $T$ will result in insufficient generation of degradation, while a high $T$ will lead to a decrease in image fidelity. Moreover, we posit that the value of $T$ should not be a manual set but rather a variable that correlates with the degradation process. In future work, we will consider proposing the auto-selection mechanism of $T$ that is specific to the degradation characteristics of given real LR and HR images, enhancing the controllability and fidelity of the model.
@@
Real-ESRGAN [33] 22.7402 0.6939
BSRGAN [32] 22.6564 0.6779
SynDiff [46] 22.4607 0.6704
@@ 4 Conclusion
Limitation: Due to the high stochasticity in the diffusion model, RealDGen sometimes finds it difficult to preserve tiny textures. In future work, we will incorporate perceptual loss for training and propose the auto-selection mechanism of the denoise step to perceive different given degradation images for further improving controllability and fidelity.
@@ Appendix A Appendix
A.1 Training detail of Content and Degradation Extractor
we initially contemplate leveraging the auto-regressive architecture of VQGAN [65] as our Content Extractor $E_{cont}$ to capture content representations. However, our experimental endeavors reveal that the Generative Adversarial Network (GAN) in VQGAN impedes fine-tuning and the extraction of realistic content representations. Consequently, we elect to employ the encoder component of VQVAE [66] as $E_{cont}$. Given that these extractors have not been trained on degradation scenarios, they are not immediately suitable for our method. To surmount this challenge, we have meticulously crafted a fine-tuning strategy predicated on reconstruction learning. Specifically, we adopt Real-ESRGAN to degrade high-resolution (HR) images to generate paired datasets for training. During training, we maintain the decoder of VQVAE in a frozen state while fine-tuning the encoder, using HR images for supervised learning, with the objective loss function being the L2 Loss. Our fine-tuning regimen is conducted with a learning rate of $1\times 10^{-5}$, utilizing 8 NVIDIA V100 GPUs, which spent approximately one week for training. The reconstructed samples rendered by our fine-tuned Content Extractor are delineated in Fig. 6. It is evident that after fine-tuning, VQVAE adeptly captures content representations in degraded scenarios and adeptly reconstructs high-resolution images.
@@ A.1 Training detail of Content and Degradation Extractor
We introduce a novel Degradation Extractor denoted as $E_{deg}$, which is comprised of a feature extraction layer integrated with 16 residual blocks and a mapping function facilitated by adaptive polling and a 4-layer convolutional structure, as shown in Fig. 7. Specifically, given the input of a low-resolution image into the Degradation Extractor, it engenders the output degradation representation $F_{deg}\in\mathbb{R}^{1\times 1\times 2048}$. To ensure the extraction of comprehensive and unique degradation representations in the LR image, we employ both reconstruction learning and contrastive learning methodologies in training our network. The training is executed on 8 NVIDIA V100 GPUs over the course of approximately seven days, with the learning rate configured at $41e-$ 4.
@@ A.3 Training Details of Decouple DDPM
We have amassed a collection of approximately 152,000 real low-resolution images sourced from publicly available datasets such as [31, 28, 49] and those captured using smartphones. To elaborate, we have extracted all low-resolution images from these datasets and have cropped each to a uniform size of $256\times 256$, resulting in a total of 110,000 images. Subsequently, to enhance the diversity of real-world degradation distribution, we have additionally procured 42,000 low-resolution images, each of the same $256\times 256$ size. Collectively, this corpus of 152,000 low-resolution images serves as the training data for our Decoupled DDPM.
@@ A.4 Detail of Modulation Block
$\mathbf{c}=\mathcal{M}\left(E_{{deg}}\left(\mathbf{x}_{{lr}}\right),E_{{cont}}\left(\mathbf{x}_{{hr}}\right)\right)$ (8)
In this section, we proceed to elucidate the intricate mechanisms underpinning the modulation block $\mathcal{M}$. $\mathcal{M}$ is meticulously constructed from a series of four modulation layers, each comprising a convolutional layer, an activation function, and a modulation unit, as shown in Fig. 8. The inputs to $\mathcal{M}$ encompass both the degradation and content representations. Within each modulation unit, the degradation representation is subjected to a sophisticated fusion process. Ultimately, the block culminates in its output through a final convolutional operation, synthesizing the enhanced representation.
@@ A.6 Stochasticity of Diffusion
As discussed in Sec. 4, to mitigate the stochasticity of diffusion, we propose a filtering mechanism to eliminate outlier data. To validate this, we conduct experiments on the RealSR dataset, calculating the similarity and data distribution compared to the real LR images in RealSR. Additionally, we verify the effectiveness using the downstream SR network SwinIR-L, with results presented in Tables 7 and 7. It can be observed that the data generated without the filtering mechanism exhibits lower similarity and further divergence from the real LR images, as well as reduced performance in the downstream SR network. This confirms the effectiveness and practicality of our proposed filtering mechanism in real-world scenarios.
@@
A.7 More Visual Comparison

### v4

@@ 3.1 Experiments Settings
Evaluation. We utilize two public benchmarks to evaluate the performance of real-world image super-resolution methods, including the real-world dataset RealSR (Cai et al., 2019) and DRealSR (Wei et al., 2020) captured by digital single-lens reflex cameras (DSLRs). To further improve the diversity and quantity of real degradation scenarios, we have collected 891 pairs of data captured by smartphones for evaluation, denoted as SmartPhone.
3.2 Quantitative Results
Generalization ability of Real SR models. We compare our method with the existing data generation methods on both PSNR-oriented and Perceptual-oriented Real SR models to validate the superiority of our method in boosting the generalization capabilities for real-world image super-resolution, and the results are shown in Table 1. We can observe that our method comprehensively improves the performance of PSNR-oriented and Perceptual-oriented SR models across three benchmarks. It’s worth noting that our approach achieves a significant performance improvement, including 0.75 dB in RRDB (Wang et al., 2018) on the PSNR of the RealSR benchmark; 0.296dB and 0.699 dB in SwinIR (Liang et al., 2021) and light-weight SwinIR-L on the PSNR of the RealSR benchmark. Furthermore, our methods also significantly improve LPIPS and FID in four SR models, including 0.0395 and 12.619 in HAT (Chen et al., 2023b) on the LPIPS and FID of the RealSR benchmark. Furthermore, we also conduct experiments on the diffusion-based model ResShift (Yue et al., 2024). Specifically, the official implementation of ResShift is trained on a large-scale dataset using Real-ESRGAN degradation with 500,000 iterations, and we perform a quick fine-tuning with 10,000 iterations using our generated data. As shown Table 2, our proposed method effectively generates accurate and realistic data in the target domain and helps ResShift quickly adapt to the new domain, yielding consistent improvements on both DRealSR and SmartPhone benchmarks. The experiments demonstrate that our method significantly improves performance across various SR approaches, including CNN-based, Transformer-based, and diffusion-based methods.
Table 2: Performance comparison of the diffusion-based ResShift model trained on our generated data versus Real-ESRGAN’s simulated data on the SmartPhone and DRealSR benchmarks.
Benchmark Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-IQA $\uparrow$
ResShift (Yue et al., 2024) 27.05 0.806 0.352 0.546
SmartPhone ResShift (Ours) 27.27 0.818 0.345 0.557
ResShift (Yue et al., 2024) 26.19 0.755 0.413 0.574
DRealSR ResShift (Ours) 26.32 0.772 0.378 0.622
@@ 3.2 Quantitative Results
Real-ESRGAN 24.3974 0.6798 0.3881 157.5237 0.2618 0.7552
BSRGAN 23.2665 0.6043 0.5093 177.9996 0.3048 0.6783
SynDiff 25.2461 0.7588 0.2371 119.9950 0.1944 0.7984
RealSR Ours 26.1615 0.7940 0.2226 107.2375 0.1865 0.8238
RealESRGAN 26.3510 0.6935 0.3842 60.2665 0.2437 0.7710
BSRGAN 26.1970 0.6869 0.4319 62.9189 0.2646 0.7137
SynDiff 28.0125 0.8136 0.1739 31.0855 0.1673 0.8462
DRealSR Ours 28.6629 0.8360 0.1497 30.0078 0.1567 0.8577
RealESRGAN 26.0680 0.5981 0.5202 106.3958 0.2855 0.7231
BSRGAN 27.6215 0.7179 0.4399 105.3289 0.2760 0.6771
SynDiff 28.5020 0.8075 0.2423 70.6379 0.2044 0.8213
@@
Accuracy of generated real LR. To validate our method’s superiority in generating accurately realistic and matched real LR, we conduct comparisons with existing methods in terms of the accuracy of the generated data on six metrics for evaluation. Specifically, we utilize HR images in the three real-world datasets as the content reference and employ the existing methods to synthesize generated LR images, while our method is able to utilize the degradation reference from real LR. As shown in Table 3, we can observe that our method achieves the best performance on three datasets under six metrics, including distribution distance metric: FID, image structure and texture similarity metric: DISTS, and perceptual metric: LPIPS and CLIP-score, etc. It demonstrates the superiority of our Decoupled DDPM and verifies that our generation method is closer to target real LR images.
Generalization Ability on Out-of-distribution Data. Although we have collected about 152,000 real LR images for training, unseen scenarios often arise in real-world applications. To further explore our generalization capabilities on out-of-distribution data, we randomly extract 200 images from the unseen real-world video SR data (Yang et al., 2021) for evaluation, denoted as RealVSR. Then, we capture the degradation representing those images to generate the training data and compare it with existing methods on the SwinIR-L model, and the results are shown in Table 6. We can observe that our method still achieves the best performance on the out-of-distribution RealVSR benchmark, surpassing the existing state-of-the-art methods by 0.2214 and 0.0055 in PSNR and SSIM, respectively. This demonstrates the ability of our method to generalize for unseen scenarios in the real world.
Figure 5: Visual comparison of generated low-resolution images with different reference real LR images. Please zoom in for better visualization.
@@ 3.3 Qualitative Results
We visualize the results of the generated LR images, as shown in Figure 3. It is clear that the Real-ESRGAN and BSRGAN, which are unable to perceive real degradation, exhibit a significant gap compared to real LR images. The SynDiff also struggles to adaptively capture the real degradation and content representation, resulting in unreal degradation distribution and low fidelity, as shown in the third row in Figure 3. However, benefiting from our adaptive perception ability of the degradation of real LR image and content of HR image, our RealDGen achieves the best visual effect in terms of both the realism of degradation and the fidelity of content. To demonstrate our method’s ability to capture the degradation distribution of arbitrary LR images, we simulate data using degradation representations from various real-world LR images and content representation from an HR image (Figure 5). In contrast to existing methods like BSRGAN and Real-ESRGAN, which rely on handcrafted degradation models and struggle to accurately reflect degradation in specific LR images, our approach separates content representation from the HR image and extracts degradation representation from the reference LR image, enabling more accurate and realistic data generation.
To further present the effectiveness of our generated data for Real SR models, we visualize the super-resolved results of the SwinIR model on the RealSR and DRealSR benchmarks, as shown in Figure 4. Models trained on SynDiff, BSRGAN, and Real-ESRGAN data exhibit noticeable artifacts and blurring, shile our method consistently delivers superior visual quality without such defects. The textual scene in the second row of Figure 4 highlights the clear advantage of our approach. Additional visual comparisons are provided in Appendix A.16.
@@ 3.4 Ablation Study
Degradation and content extractor. Our core idea is utilizing the degradation extractor $E_{deg}$ and content extractor $E_{cont}$ to extract degradation and content representation to control Decoupled DDPM in generating realistic LR. To verify this, we eliminate components $E_{deg}$ and $E_{cont}$, subsequently training the decoupled DDPM and fine-tuning the remaining extractor. For quick evaluation, we employ SwinIR-L as the baseline and evaluate on the Smartphone benchmark. The results in Table 5 demonstrate that the absence of $E_{deg}$ and $E_{cont}$ will cause mismatch and unreal degradation and content distortion problems and result in performance degradation, thereby affirming the indispensability of $E_{deg}$ and $E_{cont}$ in our method.
$T$ in data generation. As illustrated in Section 2.3, in inferencing, we denoise from an initial LR image $\mathbf{x}_{t}$ with $t$-step noise added and perform t steps of denoising to improve the fidelity of Decoupled DDPM, where $t$ is selected from the range of 0 to $T$. To explore it, we set $T$ to 200, 300, 400, and 500 and conduct ablation experiments, as presented in Table 5. It is observed that the optimal $T$ is 300. A lower $T$ will result in insufficient generation of degradation, while a high $T$ will lead to a decrease in image fidelity. We propose that the value of $T$ should be dynamically adjusted to align with the degradation level of the given LR image, rather than being manually set. In future work, we plan to develop an automatic selection mechanism for $T$ to enhance the model’s control and improve the fidelity of the generated data. More ablation studies of loss function’s hyperparameters, $n$, and $margin$, are presented in Appendix A.7.
@@
Real-ESRGAN 22.7402 0.6939
BSRGAN 22.6564 0.6779
SynDiff 22.4607 0.6704
@@ 4 Conclusion
Limitation analysis. Due to the high stochasticity inherent in diffusion models, RealDGen sometimes struggles to preserve fine textures. In future work, we plan to incorporate perceptual loss during training and develop a robust mechanism for automatically selecting the optimal denoising step based on the degradation level, thereby further enhancing the fidelity of the generated data.
Acknowledgments. This work was supported by the Natural Science Foundation of China under Grants 62225207,62436008 and 62206262.
@@ Appendix A Appendix
A.1 Training detail and analysis of Content and Degradation Extractor
A.1.1 Content Extractor
We initially contemplate leveraging the auto-regressive architecture of VQGAN Esser et al. (2021) as our Content Extractor $E_{cont}$ to capture content representations. However, our experimental endeavors reveal that the Generative Adversarial Network (GAN) in VQGAN impedes fine-tuning and the extraction of realistic content representations. Consequently, we elect to employ the encoder component of VQVAE Razavi et al. (2019) as $E_{cont}$. Given that these extractors have not been trained on degradation scenarios, they are not immediately suitable for our method. To surmount this challenge, we have meticulously crafted a fine-tuning strategy predicated on reconstruction learning. Specifically, we adopt Real-ESRGAN to degrade high-resolution (HR) images to generate paired datasets for training. During training, we maintain the decoder of VQVAE in a frozen state while fine-tuning the encoder, using HR images for supervised learning, with the objective loss function being the L2 loss. Our fine-tuning regimen is conducted with a learning rate of $1\times 10^{-5}$, utilizing 8 NVIDIA V100 GPUs, which spent approximately one week for training. The reconstructed samples rendered by our fine-tuned Content Extractor are delineated in Figure 7. It is evident that after fine-tuning, VQVAE adeptly captures content representations in degraded scenarios and adeptly reconstructs high-resolution images.
@@ A.1.1 Content Extractor
Why can the content feature extractor pre-trained on LR data robustly extract content features from HR data during data generation? The proposed content extractor is designed to robustly extract content information from an arbitrary image. Given the inevitable presence of degradation in real-world HR images, we collect 152,000 real images with various types and degrees of degradation for unsupervised fine-tuning. Within this dataset, images with relatively low degradation levels exhibit a distribution closer to that of HR images. Fine-tuning the content extractor on these real images improves its generalization capability in real-world scenarios. Quantitative results (Tables 1, 2, and 5) and qualitative results (Figures 3, 4, and 5) validate that our content extractor can effectively extract robust content representations from any HR image to generate realistic and high-fidelity LR images.
Why use LR-HR pairs to pre-train the content extractor? The content extractor is designed to extract robust ’pure’ content information from an image while ignoring potential degradation. Using LR-LR or HR-HR pairs causes the content extractor to learn identity mapping, mixing content, and degradation representations. During inference, we aim to extract content from a relatively clear image (’HR’) without involving its degradation and combine it with the degradation of another image (’LR’) to generate a new LR image. Therefore, we should use LR-HR pairs to train the content extractor. Table 2 and Figure 3 validate that our content extractor can effectively extract robust content representations to generate realistic and high-fidelity LR images.
A.1.2 Degradation Extractor
We introduce a novel Degradation Extractor denoted as $E_{deg}$, which is comprised of a feature extraction layer integrated with 16 residual blocks and a mapping function facilitated by adaptive polling and a 4-layer convolutional structure, as shown in Figure 8. Specifically, given the input of a low-resolution image into the Degradation Extractor, it engenders the output degradation representation $F_{deg}\in\mathbb{R}^{1\times 1\times 2048}$. To ensure the extraction of comprehensive and unique degradation representations in the LR image, we employ both reconstruction learning and contrastive learning methodologies in training our network. The training is executed on 8 NVIDIA V100 GPUs over the course of approximately seven days, with the learning rate configured at $41e-$ 4.
@@ A.3 Training Details of Decouple DDPM
We have amassed a collection of approximately 152,000 real low-resolution images sourced from publicly available datasets such as Wei et al. (2020); Cai et al. (2019); Ignatov et al. (2017) and those captured using smartphones. To elaborate, we have extracted all low-resolution images from these datasets and have cropped each to a uniform size of $256\times 256$, resulting in a total of 110,000 images. Subsequently, to enhance the diversity of real-world degradation distribution, we have additionally procured 42,000 low-resolution images, each of the same $256\times 256$ size. Collectively, this corpus of 152,000 low-resolution images serves as the training data for our Decoupled DDPM.
@@ A.4 Detail of Modulation Block
$\mathbf{c}=\mathcal{M}\left(E_{{deg}}\left(\mathbf{x}_{{lr}}\right),E_{{cont}}\left(\mathbf{x}_{{lr}}\right)\right)$ (8)
In this section, we proceed to elucidate the intricate mechanisms underpinning the modulation block $\mathcal{M}$. $\mathcal{M}$ is meticulously constructed from a series of four modulation layers, each comprising a convolutional layer, an activation function, and a modulation unit, as shown in Figure 9. The inputs to $\mathcal{M}$ encompass both the degradation and content representations. Within each modulation unit, the degradation representation is subjected to a sophisticated fusion process. Ultimately, the block culminates in its output through a final convolutional operation, synthesizing the enhanced representation.
@@ A.6 Stochasticity of Diffusion
As discussed in Sec. 4, to mitigate the stochasticity of diffusion, we propose a filtering mechanism to eliminate outlier data. Specifically, we extract the degradation and content representations of the three simulated LR images, compare their similarities with the input LR’s degradation and the HR’s content representations, and select the best one to enhance the fidelity and quality of the generated LR images. To validate this, we conduct experiments on the RealSR dataset, calculating the similarity and data distribution compared to the real LR images in RealSR. Additionally, we verify the effectiveness using the downstream SR network SwinIR-L, with results presented in Tables 8 and 8. It can be observed that the data generated without the filtering mechanism exhibits lower similarity and further divergence from the real LR images, as well as reduced performance in the downstream SR network. This confirms the effectiveness and practicality of our proposed filtering mechanism in real-world scenarios.
@@
A.7 More ablation studies
$n$ 1 0.1 0.01 0.001 0.0001
PSNR $\uparrow$ 28.34 28.43 28.66 28.52 28.60
FID $\downarrow$ 31.37 30.78 30.00 30.49 30.15
Table 9: Ablation studies on $n$.
$margin$ 1 2 3 4 5
PSNR $\uparrow$ 28.35 28.50 28.66 28.61 28.60
FID $\downarrow$ 30.97 30.61 30.00 30.16 30.15
Table 10: Ablation studies on $margin$.
In the loss function of the manuscript, we set the number of samples $n$ and $margin$ to 3 and 0.01, respectively, based on our experimental results. As shown in Tables 10 and 10, evaluating PSNR and FID between synthetic LR and real LR images on the DRealSR dataset, this configuration yields the best performance. Since it is impractical to exhaustively test all possible configurations, we empirically chose these values.
A.8 More downsample scale
Table 11: Comparison of data generation performance of different methods at different downsample scales on RealSR datasets.
Downsample Scale Metric BSRGAN Real-ESRGAN SynDiff Ours
$\times$4 PSNR (dB) $\uparrow$ 23.26 24.39 25.24 26.16
$\times$2 PSNR (dB) $\uparrow$ 23.26 24.39 25.24 26.16
In our manuscript, we adopt the typical downsample scale ×$\times 4$ scale setting to validate the effectiveness of the proposed data generation method. Furthermore, our method is easily adaptable to synthesizing images at different super-resolution scales, such as ×$\times 2$, and ×$\times 4$. To further demonstrate the flexibility and superiority of our approach, we present comparison results for ×$\times 4$ and ×$\times 2$. As shown in Table 11, our method achieves the best performance at both ×$\times 4$ and ×$\times 2$ scales.
A.9 Inference Time and Parameters Comparison
Table 12: Comparison of inference time with different methods.
Hand-Crafted Physical-Based Learning-Based
Methods BSRGAN Real-ESRGAN SynDiff Ours
Platform 4 CPU 4 V100 4 V100 4 V100
Times (s) 0.4267 0.1085 0.7801 0.6086
PSNR (dB) 23.26 24.39 25.24 26.16
Inference time. We test the average inference time (including IO, image processing, and generation) of producing 100 images and report the PSNR between synthetic LR and real LR on the RealSR dataset, as shown in Table 12. Using the official implementation, we measure inference time on 4 V100 GPUs. Our method adopts a sampling step of adding noise and denoising from 1 to 30, generating the highest-quality LR images and being faster than the learning-based SynDiff, although it is slower than RealESRGAN and BSRGAN. Our method and RealESRGAN can be used collaboratively and efficiently. We can use RealESRGAN for pre-training and then apply our method to create a small amount of high-quality data for fine-tuning the target scene, which does not require high speed from our method. For example, as shown in Table 2, fine-tuning ResShift on our generated realistic data significantly improves its generalization capability in real-world scenarios. Additionally, we will explore more efficient sampling strategies and distillation methods to speed up the inference of our method further.
Model parameters. Considering that BSRGAN and Real-ESRGAN are not deep-learning-based methods, following your suggestion, we compare the capacity and performance of deep learning-based Syndiff in Table 13. We see that our model not only has fewer parameters but also consistently achieves better performance on RealSR, DRealSR, and SmartPhone benchmarks.
Table 13: Parameters and performance comparison.
[…]
RealSR (Syndiff) 101.121 25.24 0.7588 0.2371 119.950 0.1944 0.7984
RealSR (Ours) 73.676 26.16 0.7940 0.2226 107.2375 0.1865 0.8238
DRealSR (Syndiff) 101.121 28.0125 0.8136 0.1739 31.0855 0.1673 0.8462
DRealSR (Ours) 73.676 28.6629 0.8360 0.1497 30.0078 0.1567 0.8577
SmartPhone (Syndiff) 101.121 28.5020 0.8075 0.2423 70.6379 0.2044 0.8213
SmartPhone (Ours) 73.676 29.0054 0.8292 0.2121 69.1162 0.1964 0.8334
A.10 Performance comparison with real data collection
Table 14: Performance comparison with existing data synthesis methods and real data collection.
[…]
RRDB (Real-ESRGAN) 24.579 0.7614
RRDB (BSRGAN) 25.406 0.7685
RRDB (SynDiff) 25.488 0.7691
RRDB (Ours) 26.238 0.7747
RRDB (Real Data) 27.302 0.7934
RRDB (Real Data+Ours) 27.302 0.7934
To further explore the performance comparison on real paired data, we further train the RRDB model using the real-collected training subset of RealSR and evaluate it on the test subset, as shown in Table 14. Since the training and test data in RealSR are captured using the same camera, they share a similar degradation distribution. Consequently, the model trained on real data achieves better performance compared to those trained on simulated data. However, collecting real-world data is often impractical, highlighting the importance of an accurate simulation method. Compared to other data synthesis approaches, our method is capable of generating realistic large-scale paired training data across diverse scenarios, enhancing the generalization capability of SR models. Furthermore, by incorporating our generated data into the real-collected training set, as shown in the last row of Table 14, the SR model’s performance improves further, demonstrating the effectiveness of our approach.
A.11 More ablation studies on degradation and content extractors
Table 15: Performance comparison before and after fine-tuning degradation and content extractor.
[…]
Real-ESRGAN 24.39 0.679 0.388
BSRGAN 23.26 0.604 0.509
SynDiff 25.24 0.758 0.237
Our methods w/o pretraining 24.13 0.761 0.301
Our methods w/o finetuning 25.93 0.786 0.234
Our methods 26.16 0.794 0.222
Table 16: Performance comparison with bicubic.
[…]
Smartphone Bicubic 28.45 0.845 0.395
Our method 28.75 0.850 0.312
RealSR Bicubic 25.98 0.755 0.442
Our method 26.23 0.774 0.297
Table 17: Performance comparison on REDS.
REDS BSRGAN Real-ESRGAN Syndiff Ours
[…]
Performance comparison before and after fine-tuning degradation and content extractor. To enhance the generalization of $E_{deg}$ and $E_{cont}$ on unseen real distributions, as shown in the main text, only a small portion of the parameters in $E_{deg}$ and $E_{cont}$ are fine-tuned. We fine-tune the last layer, as shown in Figure 2. Therefore, it’s nearly impossible for information to be directly passed through to the output. To further validate this, we conducted comparative experiments. As shown in Table 15, even without fine-tuning, our method can still generate realistic real LR, achieving the best performance compared to existing methods. After fine-tuning the extractors, the network’s generalization ability in real scenarios is further enhanced, allowing for better generation of real LR in real-world settings.
Validating the Effectiveness of Pretraining $E_{deg}$ and $E_{cont}$. To Validate the effectiveness of the pretraining stage, we removed the pretrained weights of $E_{deg}$ and $E_{cont}$ and proceeded to train them alongside DDPM. However, without pretraining, $E_{deg}$ and $E_{cont}$ were unable to distinguish between content and degradation effectively. This incapacity led to a significant decline in performance when synthesizing data, as they failed to extract degradation accurately from Real LR. Simultaneously, the content extractor struggled to capture high-fidelity content representations, resulting in generated images with lower fidelity, as shown in Table 15. We observed that without pretraining, $E_{deg}$ and $E_{cont}$ could not properly extract the respective degradations, making it challenging to maintain content consistency and degradation authenticity during the generation process, leading to a marked drop in network performance.
A.12 Comparison with Bicubic
We incorporate bicubic degradation as a baseline and compare the SR performance. To ensure a fair comparison, we utilize the officially released pre-trained RRDB models trained on bicubic data and evaluate them on the RealSR and Smartphone datasets, as shown in Tables 16. The results clearly demonstrate that the RRDB model trained on our data achieves significantly better SR performance across various evaluation metrics, particularly in the perceptual-oriented LPIPS metric. This highlights the limitations of a single bicubic model, which struggles to handle the high complexity of real-world degradations (Wang et al., 2021b; Zhang et al., 2021).
A.13 Performance evaluation on blurred LR scenarios
To evaluate performance on blurred LR scenarios, we select blurred LR images from the popular REDS benchmark (Nah et al., 2019) to test and validate the improved generalization of our proposed method in more real-world scenarios. Specifically, we evaluate RRDB models pre-trained using synthesized data from different methods. The results, as shown in Table 17, demonstrate that our method consistently achieves the best performance on out-of-distribution data with blur degradation.
A.14 Comparison with Different Contrastive Learning Methods
Table 18: Performance comparison with different contrastive learning.
[…]
DASR[D] 25.73 0.778 0.265
Our method 26.16 0.794 0.222
In (Wang et al., 2021a), the contrastive learning approach for training the degradation model constructs positive samples from different patches of the same image and negative samples from different images. However, the degradation between different patches of the same image is not always identical. For instance, in a defocused LR image, the defocused and non-defocused areas exhibit completely different degradations, and the degradation between different images may sometimes be similar. In contrast, our strategy constructs positive samples by keeping the degradation consistent while varying the content and generates negative samples by keeping the content consistent while varying the degradation, as described in Section 2.1 of the main text. This design ensures that the degradation extractor captures the complete and unique degradation distribution of the current LR image. To validate this, we replace the degradation model in (Wang et al., 2021a) and train the diffusion model, then generate real LR images on the RealSR dataset. The results, shown in Table 18, demonstrate that our proposed degradation model produces more realistic LR images compared to (Wang et al., 2021a).
A.15 More Visual Comparison on Internet real LR images
To evaluate real LR images from the internet, we compared existing methods on the SwinIR model and visualized the results, as shown in Figure 10 in the Appendix. It can be seen that our method better generates textures, while other methods tend to be over-smooth, validating the effectiveness of our approach.
Figure 10: Visual comparison of generated HR image from real internet LR images.
A.16 More Visual Comparison